Two particles with equal charge magnitudes 2X10^-7 C but opposite signs are held 15cm apart. Calculate the magnitude and direction of the electric fields at the point midway between the charges.
1
Expert's answer
2021-10-26T17:21:23-0400

"E_1=\frac{kq_1}{d^2}"

"E_2=-\frac{kq_2}{d^2}"


"E_{net}=E_1+|-E_2|"

"q_1=q_2=q"

"E_{net}=\frac{2kq}{d^2}"


"E_{net}=\frac{2\times9\times10^9\times2\times10^{-7}}{0.075^2}=64\times10^4N/C"

Direction +q to -q